OpenStudy (anonymous):

What is the distance between points (4, −7) and (−5, 9)? Round to the nearest tenth of a unit.

OpenStudy (anonymous):

Is the answer 9.1

OpenStudy (michele_laino):

we have to apply this formula \[d = \sqrt {{{\left( {9 - \left( { - 7} \right)} \right)}^2} + {{\left( { - 5 - 4} \right)}^2}} = ...?\]

OpenStudy (anonymous):

16^2 + 9^2 = 18.4?

OpenStudy (michele_laino):

d=18.4, that's right!

OpenStudy (anonymous):

Thanks!

OpenStudy (michele_laino):

Thanks!